OpenStudy (anonymous):

Please help me!!!! 1. a dartboard consists of two concentric circles the probability of hitting the inner circle is 64% it the outer circle has a diameter of 20 inches what is the radius of the inner cirlce? A. 2 inches B. 4 inches C. 8 inches D. 16 inches 2. using a sixteen-sided number cube what is the probability that you will roll an even number of an odd prime number? the number 1 isnt an odd prime. round to three decimals. A. 0.063 B. 0.813 C. 0.219 D. 0.156

OpenStudy (nirmalnema):

1) area of outer circle =3.14*400 64% of this area is the area of smaller circle. x / 3.14*400 * 100 =64 x=803.16 area of smaller circle = 3.14 *r^2 = 803.16 r^2 = 256 r = 16

OpenStudy (anonymous):

oh okay i get that thank you, i just really need help with # 2